what is the formula for the following question please:
going to the beach 5km away,Jill rides her bike at 12 km/hr,Jack jogs at 8 km/hr,how much of a head start would Jill have to give Jack to reach the beach at the same time?
thank you.

Answer by mananth(16946)

cycling speed = 12 km/h
jogging speed = 8km/h
distance = 5 km
difference in time taken =
5/8 - 5/12
LCD =24
(15-10)/24 = 3/24 = 1/8 hours more
In 1/8 hours he will jog 1/8 * 8 = 1 km
Head start = 1km